OpenStudy (anonymous):

Solve. y + 16 = –2 A. 18 B. 14 C. –14 D. –18

OpenStudy (projects):

What do you believe the answer is?

OpenStudy (anonymous):

i honestly truely dont know and i do not want to guess please help me figure out how to get to the answer?

OpenStudy (projects):

Lol, I understand, worries, but let see. You're solving for Y so go the your choices and see which one gives you -2 then comment back your answer. c:

OpenStudy (anonymous):

okay

OpenStudy (projects):

no worries** go through** sorry I'm half asleep, lol

OpenStudy (anonymous):

Let's solve your equation step-by-step. y+16=−2 Subtract 16 from both sides. y+16−16=−2−16 y=−18 Answer: y=−18

OpenStudy (anonymous):

^^

OpenStudy (projects):

Well there's your answer.. lol ^^^

OpenStudy (anonymous):

thanks for help but i honestly did not want the answer but i really appreciate it

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

YW! :)

OpenStudy (projects):

Anytime ! :D